OpenStudy (anonymous):

solve by completing the square 0=6x2+5x-5

OpenStudy (anonymous):

5=6x^2+5x 5=x(6x+5) how can I go about from here?

OpenStudy (anonymous):

instead of rearranging you could have solved for x by using the quadratic equation do you know it?

OpenStudy (anonymous):

6x^2+5x-5=0?

OpenStudy (anonymous):

sorry I tried but there was nothing to factor for... usually this becomes something like (x+/- something)(x+/- something)=0

OpenStudy (anonymous):

and if its the one with the weird suare rooting 4ab something I didn't learn that, and I would appreciate it if yu can show me the formula. I was sick the day the teacher taught it, and she told me to "ask my friends" or "look it up"

OpenStudy (anonymous):

yeah I meant the weird one which you didn't learn so we can go through it together

OpenStudy (anonymous):

I would greatly appreciate that... having a test after the break and I'm kind of screwed

OpenStudy (anonymous):

ok the formula is\[\frac{ -b pm sqrt{b^{2}-4ac} }{ 2a}\]

OpenStudy (campbell_st):

well given the fact you are asked to use completing the square that's what you should do divide every term by 6 \[x^2 + \frac{5}{6}x = \frac{5}{6}\] find half the coefficient of x and then square it... add it to both sides of the equation... then you will have a perfect square on the left.

OpenStudy (anonymous):

the system is playing up on my end but you can figure out the formula on line

OpenStudy (anonymous):

so if the value of a is 6 and b is 5 and c is -5 you slot that into the formula

OpenStudy (anonymous):

whoa what's the pm? @campbell_st I wasn't told to comple te the square, but that was the only way I thought of going since I could not factor. I should go do the other thing I'm assuming

OpenStudy (anonymous):

pm is plus/minus

OpenStudy (anonymous):

oh... ok not it makes more sense.. so it's (-1+/-sqrt(1-4(8)(-5))/16

OpenStudy (anonymous):

sorry the problem was 0=6x^2+x-5

OpenStudy (anonymous):

Here @nos since @Mariam.A since you were having trouble posting it: if the equation is: \[{ax^{2}+bx+c=0}\]\[x= \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

so it goes -1+/-sqrt(-1-(4(6)(-5))/12 -1+/-sqrt(-1+120)/16 -1+/-sqrt(119)/16 and -1+/-10sqrt(19)/16 -1/16+/-5sqrt(19)/8

OpenStudy (anonymous):

more or less?

OpenStudy (anonymous):

you messed up the b^2. your first line should be: -1+/-sqrt(1-(4(6)(-5))/12

OpenStudy (anonymous):

oh so the sqrt(121) is the correct answer? then it would be 11

OpenStudy (anonymous):

-1+/-11/12 -12/12=-1 10/12=5/6

OpenStudy (anonymous):

yes. The equation you gave is actually factorable: \[0=6x^2+x-5\]\[0 = (6x-5)(x+1)\]

OpenStudy (anonymous):

<I just need to try those answers>

OpenStudy (anonymous):

wow. that lost the purpose it's ok at least it teached me how to do the weird thing, and I'm actually thankful for it.

OpenStudy (anonymous):

I should check my answers, but thanks for the steps on the equation :)

OpenStudy (anonymous):

yes that weird thingy that you said is actually very helpful so it would be wise to learn how to do it.